# Accelerated Interactive Auralization of Highly Reverberant Spaces using Graphics Hardware

### v1

@@ VII Conclusion
In this paper, we presented a real-time implementation of a GPU-accelerated loudspeaker-based auralization system for the interactive synthesis of large room acoustics with integrated feedback cancellation. The system was designed to be flexible and applicable to a wide range of loudspeaker-based auralization scenarios, with virtually no limitation on the length of the synthesis filters and the number of channels in the reproduction setup. The implementation was evaluated in terms of processing time by comparing the performance of the GPU and CPU implementations. It was shown that the GPU implementation achieves a performance gain of up to $60$ times compared to the CPU implementation, enabling real-time auralization at $48$ kHz for synthesis filters with a duration of up to $20$ seconds and $32$ output channels. The results demonstrated that the proposed implementation was capable of performing real-time auralization with minimal latency, making it suitable for interactive applications in large room acoustics. While the Python implementation of the system was not optimized for real-time performance, it was shown to achieve real-time performance using accelerated libraries. Future work will focus on further optimizing the system for real-time performance, and on interactive perceptual evaluation of the system.
VIII Acknowledgments

### v2

@@ VIII Conclusion
This paper presented a real-time implementation of a GPU-accelerated loudspeaker-based auralization system for the interactive synthesis of large room acoustics with integrated feedback cancellation. The system was designed to support scalable loudspeaker-based synthesis with practical flexibility in filter length and channel count. The implementation was evaluated in terms of processing time by comparing the performance of the GPU and CPU implementations. It was shown that the GPU implementation achieves a performance gain of up to $60$ times compared to the CPU implementation, enabling real-time auralization at $48$ kHz with a block size of $128$ samples for synthesis filters with a duration of up to $20$ seconds and $32$ output channels. This block size corresponds to a buffer latency of $2.67$ ms, and an end-to-end latency of $12.26$ ms as measured using an RME Fireface UFX+ audio interface connected to the auralization system. The results demonstrated that the implementation supports real-time auralization with minimal latency, making it suitable for interactive applications with large room acoustics. While the Python implementation of the system was not optimized for real-time performance, it was shown to achieve real-time performance using accelerated libraries. Future work will focus on further optimizing the system for real-time performance, and on interactive perceptual evaluation of the system.
IX Acknowledgments